Working with complex numbers and I need to find all the real numbers
x and y that satisfy the equation 3x + 1 + 2yi = x – 5i
Thank you.

Answer by AnlytcPhil(1806)

Working with complex numbers and I need to find all the real numbers x and y
that satisfy the equation 3x + 1 + 2yi = x – 5i
I will color all the terms which contain "i" red and those that
don't contain "i" blue.
3x + 1 + 2yi = x - 5i
Now equate only the blue terms on the left
with those on the right
3x + 1 = x
2x = -1
x = -1/2
Now equate only the red terms:

2yi = –5i
y = -5i/(2i)
y = -5/2

Edwin